if short run elasticity in 0.3 and price falls by 30% what is change in quantity ?
1
Expert's answer
2017-05-09T09:47:09-0400
If E = 0.3, dP = - 0.3P, then E = (dQ/Q)/(dP/P), dQ/Q = 0.3*(-0.3P/P) = -0.09.
So quantity falls 9 %.